OpenStudy (anonymous):

Find the limit (if it exists). lim x->0 [[1/(3+x)]-(1/3)]/x

hartnn (hartnn):

\(\huge\frac{1}{3+x}-\frac{1}{3}=\frac{3-(3+x)}{3(3+x)}=??\)

hartnn (hartnn):

can u solve further?

OpenStudy (anonymous):

yes, thank you!

hartnn (hartnn):

welcome :) let me know your final answer,so that we can verify.

OpenStudy (anonymous):

I get 0 but my book says -1/9?

hartnn (hartnn):

u have: \(\huge\frac{1}{3+x}-\frac{1}{3}=\frac{3-(3+x)}{3(3+x)}=\frac{-x}{3x(3+x)}\) right ?

OpenStudy (anonymous):

No, I thought I could just solve by substituting 0 for x.

OpenStudy (anonymous):

I dont think you notcied that the orignal equation was divided by x.

hartnn (hartnn):

but u have x in denominator also , right ? so u just cannot put x=0 cancel out x from numerator and denominator then put x=0

OpenStudy (anonymous):

Oh, i see now. thank you.

hartnn (hartnn):

did u get -1/9 ?

OpenStudy (anonymous):

Yes. But could you please explain how you got from 3-(3+x)/3(3+x) to -x/3x(3+x)? I was a bit confused about that.

hartnn (hartnn):

the numerator, right ? 3-(3+x) = 3-3-x = 0-x = -x ok? denominator is same.

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome :)